# Talent Hoarding in Organizations

### v2

@@ 3.3 Survey of Managers
In addition to the employee survey that was discussed in Section 2.1, I conduct a survey of managers at the firm that was designed to elicit the hoarding incentives managers face and the actions that may follow them. Managers in the firm were invited via e-mail by the firm’s HR department. Respondents were prompted to provide their perspectives on managers’ role in developing talent at the firm based on their own experiences as managers. The survey received responses from 62% of managers, yielding over 3,000 responses. Appendix Table D2 shows that managers who responded to the survey are similar in their demographics to managers in my analysis sample. The primary objective of the survey was to directly elicit the extent to which managers perceive misaligned incentives with respect to the development of talent in the firm. Managers were asked to respond to a series of questions including explicit statements with respect to potential talent hoarding. The survey also carries the advantage that it does not rely on worker reports, allowing me to more directly infer the magnitude of managerial talent hoarding in the firm. See Appendix Section D.3 for an abbreviated version of the survey instrument.
@@ 4 Evidence from Managers: Hoarding Incentives and Actions
The conceptual framework in Section 2.2 predicts that managers should hoard workers on their team due to the misaligned incentives they face. This section uses a survey of managers in the firm to provide direct evidence that managers perceive misaligned incentives that lead to talent hoarding. I then leverage a direct measure of managers’ hoarding actions using the firm’s HR data to show that managers who face stronger hoarding incentives are more likely to take hoarding actions.
@@ 4.1 Manager-Reported Hoarding Incentives and Actions
Responses to the survey of managers described in Section 3.3 reveal that managers face strong incentives to hoard talent instead of supporting workers in their career development outside the team. The majority (55%) of managers explicitly acknowledge that supporting workers’ career development entails a conflict of interest because more developed workers are more likely to leave the team (Figure 1). This conflict of interest appears to be driven by the fact that managerial talent development is not sufficiently rewarded by the firm: only 36% of managers perceive their impact on workers’ careers to be valued by the firm at least as much as their impact on immediate team performance. Yet managers can have important impacts on their subordinates’ careers: 96% of managers believe that their direct intervention, for instance through encouraging workers or increasing their visibility outside the team, has a large impact on their workers’ career development. Although the broad majority of managers (87%) indicate that talent development is an area where managers can create a high impact for the firm, only 40% report that a record of success with respect to their workers’ career development is important for their compensation and promotion prospects. This difference indicates that although managers could benefit the firm by supporting their workers’ career development, managers perceive a lack of incentive to do so. [^13] See Appendix Table B2 for survey quotations of managers describing this lack of incentive. A key disincentive to develop worker talent is that developing workers increases the risk of them leaving the team. When asked under which conditions managers would be more likely to support workers’ career development, the most common response among managers (68%) is if replacing employees who leave the team would be easier (Panel A of Appendix Figure A1). In addition, 42% of managers indicate that being able to develop workers within the team, instead of losing them to other teams, would make them more likely to engage in talent development. These findings provide direct evidence that the risk of losing team members deters managers from supporting their workers’ career development. Managers also indicate that increasing the returns to talent development would make them more likely to support workers’ careers. Most importantly, 52% of managers state receiving more recognition for developing talent as key condition. Direct impacts on managers’ promotion prospects and financial incentives, which are not common features of the internal labor market I study, were only named by 22% and 21% of managers, respectively. Managers’ survey responses indicate that managers have ample scope to impact their workers’ career development. Panel B of Appendix Figure A1 shows that when asked about specific actions they can take, managers most frequently cite making workers visible outside the team (84%) and establishing an open dialogue with workers about career development (84%). Managers also identify allowing workers to participate in trainings (69%), encouraging employees to pursue different job opportunities (62%), providing guidance for career planning (58%), and notifying workers about job opportunities outside the team (32%). These actions also accord with the survey responses provided by workers (discussed in Section 2.1), who report that talent hoarding typically occurs through managers withholding career-planning support, discouraging workers from applying to internal positions, and reducing the visibility of worker ability (e.g., by suppressing public signals of worker ability or restricting access to high-visibility trainings). Workers’ survey responses also suggest that managerial talent hoarding actions are effective in reducing workers’ career opportunities. For example, workers who report talent hoarding in form of retaliation are 5.1 percentage points (56%) less likely to have received a job recommendation by their manager ($p$=0.000), and 4.6 percentage points (13%) less likely to have received job recommendations from anyone in the firm ($p$=0.000). Together, these results point to several complementary ways in which managers can effectively impact workers’ career development. How widespread is talent hoarding? I use two complementary survey questions to assess the magnitude of managerial talent hoarding at the firm. First, my preferred measure is based on the question, “How often might leaders find themselves in situations where they need to dissuade a team member from exploring opportunities in another department due to immediate team needs or performance goals?”. The phrasing of this question was specifically designed to reflect the definition of talent hoarding in the conceptual framework: managers deliberately reducing the likelihood that a worker leaves the team. I find that 75% of managers respond that this situation occurs at least sometimes, indicating that the broad majority of managers explicitly recognize the existence of talent hoarding. My second measure is based on asking managers to report “the reasons that may prevent leaders from investing time and effort towards their employees’ career development.” Note that this question is more indirect because it asks about foregone investment rather than actively dissuading worker mobility. Nonetheless, I find that 45% of managers cite the risk of losing workers as a key reason for not investing in employees. Moreover, 66% of managers cite the need to prioritize short-term targets over long-term employee development, which accords with the responses to the first question and underscores the importance of team performance in managerial incentives. Collectively, these responses corroborate the finding that talent hoarding is a common reaction to prevailing managerial incentives.
@@ 4.2 Measuring Hoarding Actions in the Data
Measuring managers’ talent hoarding actions in the data is difficult because it typically occurs through interpersonal interactions. The survey evidence suggests that one common way in which managers limit workers’ opportunities to leave the team is by suppressing worker visibility. To capture high-stakes talent hoarding actions in the data, I identify managers’ systematic suppression of worker visibility based on a measure of worker visibility that I collect from the firm’s HR databases. Specifically, I measure systematic differences between potential ratings and performance ratings. Performance ratings are meant to provide task-specific feedback to workers about their past performance in their current position. Potential ratings are designed to inform the firm about a worker’s future potential for higher-level positions and thus are meant to identify workers who would be likely to perform well if they were promoted. Both ratings are conducted simultaneously by a worker’s direct supervisor and are very similar to common worker assessments, such as the nine-box grid, that are used by many organizations across the world (Cappelli and Keller, 2014). An important distinction between performance and potential ratings is the extent of their visibility outside of a worker’s team. Performance ratings are private signals to the worker and are not shared with other units in the firm. If a worker applies for a job in a different unit, that unit will not be able to access the worker’s past performance ratings. The firm’s HR department describes the privacy norms regarding performance ratings as similar to the strong taboo in Germany of talking about one’s salary. In contrast, potential ratings are public signals of worker talent and are widely circulated within the firm. The HR department regularly circulates lists of high-potential workers, making them highly visible outside of worker’s current team. Intuitively, a manager who wants to hoard talent should give workers lower (public) potential ratings relative to their (private) performance ratings. A survey of the firm’s employees suggests that this method of talent hoarding is commonplace. [^14] For instance, one worker states, “Supervisors suppress potential ratings because of fear that employees will leave their current position for a promotion.” (Appendix Table B1). For a given worker, the difference between performance and potential ratings may reflect worker-specific factors that are unrelated to talent hoarding; however, because managers have discretion in conducting these evaluations, comparing systematic differences between the ratings across workers can capture manager behavior. [^15] Illustrating the importance of manager discretion in this setting, Benson et al. (2021) find in contemporaneous work that managers in a large retailer overrate men’s potential compared to women’s, possibly as a reaction to men’s higher turnover risk. My measure of talent hoarding is defined by the difference between the actual and predicted potential ratings a manager assigns to their workers. Specifically, I take the residuals from the OLS estimation of the following regression of the potential rating given to worker $i$ in quarter $t$ by manager $m$ on their performance rating and other worker characteristics:
@@
In Equation 2, $X_{it}$ denotes the vector of controls described in Section 3.2, $\beta_{t}$ represents quarter fixed effects. I compute the average difference (over workers and quarters) between predicted and actual potential ratings for each manager. I classify a manager as having a high propensity to hoard talent if this difference is in the top tercile of the manager distribution, indicating that the potential ratings the manager assigns are systematically lower than what one would expect. Appendix Section E provides additional information on the construction and validity of this measure. I find that the measure is reasonably stable over time, supporting the systemic notion of talent hoarding the measure is meant to capture. The talent hoarding measure is also not simply a function of managers’ ability to assess talent, suggesting that differences between performance and potential ratings are unlikely to be a result of managers’ involuntary mistakes. In addition, the underrating of potential captured by this measure does not appear to represent managers’ correct assessment of workers’ future performance. Instead, managers’ suppression of the public signals of worker talent is consistent with their incentive to reduce visibility outside of the team. This measure allows me to directly test empirical predictions for the circumstances under which managers should engage in more hoarding (Section 4.3) as well as to document how talent hoarding depresses workers’ career development (Section 5).
@@ 4.3 Testing Empirical Predictions for Managerial Hoarding Actions
To provide evidence that managers purposefully hoard talent based on managers’ revealed actions, I use the direct measure of talent hoarding I construct to assess whether hoarding actions are more likely to occur when managers face greater hoarding incentives. Following the empirical predictions from my conceptual framework in Section 2.2, I relate my measure of managerial talent hoarding for each of the over 6,000 managers in my sample to the incentives these managers face. Specifically, I analyze whether managers are more likely to be classified as hoarding-prone based on the measure discussed in Section 4.2 if they face stronger incentives to hoard. To conduct this test, I estimate a linear model for managers’ hoarding propensity using an OLS regression on an indicator that managers face a strong incentive to hoard and a set of manager controls (i.e., gender, German citizenship, age, firm tenure, and education). Motivated by the conceptual framework, I construct three proxy measures that capture variation in managers’ hoarding incentives. First, the framework predicts that managers should engage in more talent hoarding if the gains from hoarding are larger. To proxy for these gains, I classify managers based on the share of their overall compensation that is based on team performance. Managers who I classify as having high financial incentives receive three times higher variable compensation relative to those with low financial incentives, implying that losing a talented worker is more financially costly to them. Second, the framework predicts that talent hoarding should be greater if a worker’s departure has a larger impact on team performance. I proxy for this dimension using the size of the manager’s team. Intuitively, smaller teams have fewer people who can compensate for a single worker’s departure. Managers are classified as having strong hoarding incentives if their team has four or less teammates, corresponding to the bottom quartile of team size at the firm. Third, the framework predicts more hoarding if managers incur lower costs of talent hoarding. I proxy for this dimension using variation in the visibility of worker talent. In functional areas where individual achievements are difficult to observe (e.g., corporate functions such as HR or marketing), it should be very easy for managers to hoard talent because the manager has a high degree of control over worker visibility. On the other hand, in areas such as engineering where many workers file patents, it should be more difficult for managers to hoard. [^16] High-visibility functions are defined as engineering, project management, and quality management. Low-visibility functions are defined as administration, finance, safety and health, HR, marketing, it, logistics, and purchasing. Managers who face greater hoarding incentives are substantially more likely to engage in talent hoarding. Panel A of Figure 2 plots the share of managers that are classified as being hoarding-prone based on my measure of suppressing potential ratings, separately by proxies of hoarding incentives. Managers who face high financial incentives are 31% more likely to hoard talent ($p$=0.000). When using team size to proxy for hoarding incentives, I find that managers in small teams are 38% more likely to engage in talent hoarding ($p$=0.000). Finally, managers in functions with low talent visibility are 13% more likely to hoard ($p$=0.002). These results show that across each of the three proxies for hoarding incentives, hoarding actions are much more common among managers who face stronger incentives to hoard, confirming the predictions from my conceptual framework. Higher levels of talent hoarding for managers facing stronger incentives are also visible in several complementary measures of hoarding actions, which corroborate the result based on potential ratings. First, I follow survey responses indicating that restricting access to high-visibility trainings is another common way through which managers are able to suppress worker visibility. In my sample, 2% of workers are assigned by their manager in a given quarter to participate in a high-visibility training. These are prestigious in-person trainings that are organized by a centralized and well-recognized department at the firm, thereby increasing worker visibility outside the team. For example, lists of training participants are widely circulated within the firm, and many workers list trainings on their CVs. I use information on which workers a manager assigns to these high-visibility trainings to construct a complementary measure of managers’ hoarding propensity. To do so, I follow the approach in Section 4.2 and identify hoarding-prone managers based on the mean deviation between expected and realized training assignments, replacing potential ratings with training assignments in Equation 2. Similar to the results using my primary measure of talent hoarding, I find that managers are more likely to engage in hoarding in the form of restricting access to high-visibility trainings if they face greater hoarding incentives. Panel B of Figure 2 shows that managers with high financial incentives to hoard are 53% more likely to engage in talent hoarding ($p$=0.000). With respect to team size, a similar pattern emerges with managers in small teams being 26% more likely to hoard ($p$=0.000). Managers who work in areas with low talent visibility are 10% more likely to hoard ($p$=0.021). Overall, these patterns are nearly identical to those using the measure of hoarding based on potential ratings. A second source of corroborating evidence comes from managers’ survey responses. The advantage of these survey-based measures is that they identify talent hoardings based on managers’ explicit statements. Panel A of Table 2 presents the share of managers who report that having to dissuade team members from exploring opportunities outside the team is a situation that occurs at least sometimes. Column 1 shows that managers with greater financial incentives are 11% more likely to report talent hoarding ($p$=0.017). Column 2 shows that managers who work in small teams are 12% more likely to report hoarding ($p$=0.000), while those that face lower hoarding costs because they work in an area with low talent visibility are 9% more likely to report talent hoarding ($p=$0.277, Column 3). Panel B of Table 2 documents similar patterns based on the share of managers who cite the risk of losing workers as a reason why managers refrain from supporting workers’ career development. Managers facing greater financial incentives are 19% more likely to report talent hoarding ($p$=0.070), while those in smaller teams are 17% more likely to do so ($p$=0.002). Managers in areas with low talent visibility are 32% more likely to report hoarding ($p$=0.010). Taken together, these results provide evidence using several complementary measures of manager actions that managers are more likely to engage in talent hoarding if they face greater hoarding incentives. Interpreted through the lens of my conceptual framework, these results confirm that more misaligned incentives lead to higher levels of talent hoarding.

### v3

@@ 3.4 Survey of Managers
In addition to the employee survey that elicited worker perceptions, I conducted a survey of managers at the firm. The primary objective of the survey was to directly elicit the extent to which managers perceive misaligned incentives with respect to the development of talent in the firm. This survey offers the advantage that it does not rely on worker reports, allowing me to more directly elicit the extent of managerial talent hoarding in the firm.
Following the firm’s standard survey protocol, managers in the firm were invited to participate in the survey via e-mail by the firm’s HR department. Respondents were prompted to provide their perspectives on managers’ role in developing talent at the firm based on their own experiences as managers. The survey received responses from 62% of managers, yielding over 3,000 responses. Managers were asked to respond to a series of questions including explicit statements with respect to potential talent hoarding. See Appendix Section D.3 for the survey instrument and the distribution of the responses.
Several features of the survey were designed to foster trust in the survey instrument, and to frame the survey as an opportunity to provide feedback to the firm’s HR department. First, the survey invitation emphasized anonymity, explicitly assuring that responses could not be linked to individual managers (Appendix Figure D2). Rather, it framed the survey as an opportunity for managers to provide confidential input on improving HR practices. Additionally, the invitation noted that the survey had been approved by the firm’s workers’ council, signaling that it had been vetted to protect participants’ confidentiality. Second, before the survey launch, HR employees were briefed and instructed to address manager inquiries, reinforcing the survey’s legitimacy and approval by the firm. Third, the firm’s leadership council, a prominent and influential internal body advocating for managers’ interests, publicly endorsed the survey and encouraged managers to participate. The high response rate of 62% of managers at the firm indicates strong engagement and suggests that managers responded favorably to these design features. Moreover, the fact that the majority of respondents (62%) voluntarily provided written suggestions on how the firm could better support leaders suggests that managers saw the survey as a meaningful channel for providing anonymous feedback.
The fact that the survey invitation explicitly stated that participation was anonymous and could neither be linked to individual managers and was not permitted to be used for monitoring, suggests that respondents’ concerns about reputational risks of admitting to talent hoarding in the survey were small. A related concern is that social desirability bias may lead managers to underreport talent hoarding because it is perceived as detrimental to the firm. To minimize such bias, the survey invitation and instrument (Appendix D.3.1) both avoided emphasizing talent hoarding, and the bulk of questions elicit managers’ perceptions of their firm’s general policies and leadership behaviors, rather than managers’ own actions. In light of the negative view of talent hoarding, any remaining social desirability bias would likely lead to an underestimation of its prevalence (see Appendix Section D.3.4 for a more detailed discussion).
A final concern regarding the interpretation of the survey results is that respondents may be a selected sample of all managers. Since the survey invitation did not explicitly mention talent hoarding, it is unlikely that managers deliberately sorted into the survey by their propensity to hoard talent. An analysis of respondent characteristics shows that survey participants are largely representative of the overall sample in terms of their demographics, experience, and functional area (Appendix Table D2). In addition, when comparing managers who responded before a reminder was sent to those who responded after (likely more reluctant participants), I find that the demographics of early and late respondents are similar, further suggesting that non-random selection in the survey is unlikely to meaningfully affect my results (Appendix Table D3).
@@ 4 Evidence from Managers: Hoarding Incentives and Actions
The conceptual framework in Section 2.2 predicts that managers should hoard workers on their team when they face incentives to retain talent that conflict with broader organizational goals. In this setting, such misaligned incentives are likely to be particularly severe when a manager’s compensation depends heavily on team performance, when a worker’s departure would be particularly costly (e.g., in small teams), or when hoarding talent is less costly because individual contributions are hard to observe (e.g., in low-visibility functions). This section uses a survey of managers in the firm to provide direct evidence that managers perceive misaligned incentives that lead to talent hoarding. I then leverage a direct measure of managers’ hoarding actions using the firm’s HR data to show that managers who face stronger hoarding incentives are more likely to take hoarding actions.
@@ 4.1 Manager-Reported Hoarding Incentives and Actions
Responses to the manager survey described in Section 3.4 reveal that managers face strong incentives to hoard talent instead of supporting workers in their career development outside the team.
To assess whether managers perceive a tension between developing talent and retaining employees, the manager survey asked respondents to rate their agreement with the following statement: “Talent development entails a conflict of interest for leaders, because more developed employees are more likely to leave the team.” Managers responded on a four-point scale ranging from “I strongly agree” to “I totally do not agree.” See Appendix Table D9 for the raw response distribution for this statement. The majority (55%) of managers agree or strongly agree, thereby explicitly acknowledging that supporting workers’ career development entails a conflict of interest because more developed workers are more likely to leave the team (Bar 1 of Figure 1). [^13] Note that to the extent that social desirability bias affected manager responses, this figure likely represents an underestimate of the share of managers who perceive a conflict of interest.
To better understand whether the perceived conflict of interest reflects a lack of rewards for managers’ talent development efforts, the manager survey included two additional statements rated on the same four-point agreement scale. First, to capture how talent development is valued relative to a core managerial responsibility—team performance—respondents were presented with the statement: “The firm values a manager’s impact on employee career development at least as much as their impact on team performance.” Second, respondents were asked to indicate agreement with the statement: “A leader’s direct intervention (e.g., encouragement, increasing visibility) has a large impact on employees’ career development.” This second question serves as a useful comparison, as it contrasts how managers perceive the rewards for talent development with their perceived ability to take meaningful action. See Appendix Table D9 for raw response distributions.
Responses to these questions suggest that managers perceive effort expended on talent development as having high impact but yielding limited rewards. While 96% of managers believe that their direct intervention has a large impact on their workers’ career development, only 36% of managers perceive their impact on workers’ careers to be valued by the firm at least as much as their impact on immediate team performance (Bars 2 and 3 of Figure 1).
To further probe the robustness of this finding, I draw on a separate set of survey questions that use a multiple-choice format rather than an agreement scale. These questions provide direct benchmarks of talent development relative to other core managerial responsibilities. To assess perceived impact, managers were asked: “Where do you think leaders can create the most impact for [Company]?” Responses were recorded on a four-point scale ranging from “Very high impact” to “Very low impact.” To contrast this with perceived rewards, managers were also asked: “When it comes to leaders’ compensation and promotion prospects, how important do you think is a record of success in each of the following?” This question used a similar four-point scale, ranging from “Very important” to “Not important.” Both questions asked managers to evaluate six core responsibilities that are considered central to middle management at the firm: strategic orientation, vision, operational efficiency, talent development, culture and morale, and innovation. See Appendix Tables D4 and D5 for raw response distributions.
The survey responses reveal a notable gap between perceived impact and perceived reward for talent development. While a large majority of managers (87%) report that talent development is an area where they can create high impact for the firm, only 40% indicate that a track record in developing employees is important for their own compensation and promotion prospects. This gap corroborates the finding that despite recognizing the potential benefits of investing in their employees’ careers, managers perceive a lack of incentives for doing so. [^14] See Appendix Table B2 for survey quotations of managers describing this lack of incentives. For instance, one manager stated: “Managers pursue their own goals and often prevent further development of workers, because they are not rewarded for developing talent.”
Comparing managers’ attitudes about talent development and their other core responsibilities are similarly revealing. First, when comparing the perceived importance of different responsibilities for managers’ compensation and promotion prospects, talent development emerges as the lowest-rated dimension (Panel B of Appendix Figure D3). Second, the discrepancy between perceived impact (Panel A) and perceived reward (Panel B) is more pronounced for talent development than for any other responsibility, suggesting a particularly sharp misalignment between what managers believe benefits the firm and what is recognized in their own advancement.
In the appendix, I present additional results that benchmark responses across these different survey questions. For example, when comparing these results to the earlier agreement-based questions (specifically, the share of managers who agree that talent development is valued by the firm and that their actions meaningfully influence workers’ careers), the patterns are strikingly similar. See Appendix Section D.3.4 for details on this comparison and a discussion of additional benchmarking exercises. Together, these survey findings support the interpretation that talent development is viewed as under-incentivized by managers, despite being recognized as impactful.
How widespread is talent hoarding? I use two complementary survey questions to assess the magnitude of managerial talent hoarding at the firm. First, my preferred measure uses a survey question that was designed to directly reflect the definition of talent hoarding from the conceptual framework: managers deliberately reducing the likelihood that a worker leaves the team. The question asked, “How often might leaders find themselves in situations where they need to dissuade a team member from exploring opportunities in another department due to immediate team needs or performance goals?” Responses were measured on a four-point scale ranging from “Never” to “Very often.” See Appendix Table D10 for the raw response distribution. I find that 75% of managers respond that this situation occurs at least sometimes, indicating that the broad majority of managers explicitly recognize the existence of talent hoarding.
My second measure probes the issue more indirectly by focusing on foregone investment in employee development rather than actively dissuading worker mobility. Managers were asked, “What are the reasons that may prevent leaders from investing time and effort towards their employees’ career development?” Respondents could select multiple response options among the following: leaders’ lack of knowledge, limited resources, risk of losing talent, need to prioritize short-term targets, and lack of interest on the part of the employee. See Appendix Table D7 for the raw response distribution. I find that 45% of managers cite the risk of losing workers as a key reason for not investing in employees. Moreover, 66% of managers cite the need to prioritize short-term targets over long-term employee development, which accords with the responses to the first question and underscores the importance of team performance in managerial incentives. Collectively, these responses corroborate the finding that talent hoarding is a common reaction to prevailing managerial incentives.
@@ 4.2 Measuring Hoarding Actions in the Data
Measuring managers’ talent hoarding actions in the data is difficult because it typically occurs through interpersonal interactions. The survey evidence suggests that one common way in which managers limit workers’ opportunities to leave the team is by suppressing worker visibility. To capture high-stakes talent hoarding actions in the data, I identify managers’ systematic suppression of worker visibility based on a measure of worker visibility that I collect from the firm’s HR databases. Specifically, I measure systematic differences between potential ratings and performance ratings. Performance ratings are meant to provide task-specific feedback to workers about their past performance in their current position. Potential ratings are designed to inform the firm about a worker’s future potential for higher-level positions and thus are meant to identify workers who would be likely to perform well if they were promoted. Both ratings are conducted simultaneously by a worker’s direct supervisor and are very similar to common worker assessments, such as the nine-box grid, that are used by many organizations across the world (Cappelli and Keller, 2014).
An important distinction between performance and potential ratings is the extent of their visibility outside of a worker’s team. Performance ratings are private signals to the worker and are not shared with other units in the firm. If a worker applies for a job in a different unit, that unit will not be able to access the worker’s past performance ratings. The firm’s HR department describes the privacy norms regarding performance ratings as similar to the strong taboo in Germany of talking about one’s salary. In contrast, potential ratings are public signals of worker talent and are widely circulated within the firm. The HR department regularly circulates lists of high-potential workers, making them highly visible outside of worker’s current team.
Intuitively, a manager who wants to hoard talent should give workers lower (public) potential ratings relative to their (private) performance ratings. A survey of the firm’s employees suggests that this method of talent hoarding is commonplace. [^15] For instance, one worker states, “Supervisors suppress potential ratings because of fear that employees will leave their current position for a promotion.” (Appendix Table B1). For a given worker, the difference between performance and potential ratings may reflect worker-specific factors that are unrelated to talent hoarding; however, because managers have discretion in conducting these evaluations, comparing systematic differences between the ratings across workers can capture manager behavior. [^16] Illustrating the importance of manager discretion in this setting, Benson et al. (2021) find in contemporaneous work that managers in a large retailer overrate men’s potential compared to women’s, possibly as a reaction to men’s higher turnover risk.
My measure of talent hoarding is defined by the difference between the actual and predicted potential ratings a manager assigns to their workers. Specifically, I take the residuals from the OLS estimation of the following regression of the potential rating given to worker $i$ in quarter $t$ by manager $m$ on their performance rating and other worker characteristics:
@@
In Equation 6, $X_{it}$ denotes the vector of controls described in Section 3.2, $\beta_{t}$ represents quarter fixed effects. I compute the average difference (over workers and quarters) between predicted and actual potential ratings for each manager. This defines a manager-level measure of talent hoarding, reflecting a manager’s systematic tendency to assign public potential ratings that are higher or lower than expected given worker performance and characteristics based on all workers in their team during my sample period.
Figure 2 shows the distribution of the mean deviation between predicted and actual potential ratings. The data for this figure is at the manager level (i.e., containing one observation for each of the over 6,000 unique managers in my sample). The distribution is centered close to zero (median = –0.006, SD = 0.212) and reveals substantial variation. Many managers assign potential ratings roughly in line with predictions, but there is meaningful mass in both tails. Managers in the left tail appear to inflate worker visibility by assigning higher-than-expected potential ratings. In contrast, managers in the right tail consistently assign lower-than-expected potential ratings, suppressing worker visibility in a manner consistent with strategic talent hoarding.
In my empirical analysis, I use both the continuous measure of a manager’s mean deviation between predicted and actual potential ratings, and discrete classifications based on terciles of this distribution. The continuous measure enables a graded interpretation: larger positive deviations correspond to greater suppression of worker visibility. To analyze meaningful variation in managerial behavior, I classify managers in the top tercile (mean deviation above 0.1036, shaded in orange) as having a higher tendency to suppress visibility. This classification allows me to focus on managers who are most likely engaged in systematic hoarding behavior, and thus also allows a direct comparison with the survey-based results that rely on binary measures of managerial hoarding actions. Robustness tests in Appendix Sections E.2 and E.3 confirm that results are consistent when using the continuous measure, a median split, or a more conservative quantile cutoff.
Appendix Section E.1 provides additional information on the construction and validity of this measure. I show that manager-level hoarding measure are stable over time and not simply driven by noise or measurement error. The talent hoarding measure is also not simply a function of managers’ ability to assess talent, suggesting that differences between performance and potential ratings are unlikely to be a result of managers’ involuntary mistakes. In addition, the underrating of potential captured by this measure does not appear to represent managers’ correct assessment of workers’ future performance. Instead, managers’ suppression of the public signals of worker talent is consistent with their incentive to reduce visibility outside of the team.
This measure allows me to directly test empirical predictions for the circumstances under which managers should engage in more hoarding (Section 4.3) as well as to document how talent hoarding depresses workers’ career development (Section 5).
@@ 4.3 Testing Empirical Predictions for Managerial Hoarding Actions
To provide evidence that managers purposefully hoard talent based on managers’ revealed actions, I use the direct measure of talent hoarding I construct to assess whether hoarding actions are more likely to occur when managers face greater hoarding incentives. Following the empirical predictions from my conceptual framework in Section 2.2, I relate my measure of managerial talent hoarding for each of the over 6,000 managers in my sample to the incentives these managers face. Specifically, I analyze whether managers are more likely to be classified as hoarding-prone based on the measure discussed in Section 4.2 if they face stronger incentives to hoard. To conduct this test, I estimate a linear model for managers’ hoarding propensity using an OLS regression on an indicator that managers face a strong incentive to hoard and a set of manager controls (i.e., gender, German citizenship, age, firm tenure, and education).
Motivated by the conceptual framework, I construct three proxy measures that capture variation in managers’ hoarding incentives. First, the framework predicts that managers should engage in more talent hoarding if the gains from hoarding are larger. To proxy for these gains, I use information from managers’ variable compensation to measure the portion of their overall compensation that is related to team performance. Second, the framework predicts that talent hoarding should be greater if a worker’s departure has a larger impact on team performance. I proxy for this dimension using the size of the manager’s team. Intuitively, smaller teams have fewer people who can compensate for a single worker’s departure. Third, the framework predicts more hoarding if managers incur lower costs of talent hoarding. I proxy for this dimension using variation in the visibility of worker talent. In functional areas where individual achievements are difficult to observe (e.g., corporate functions such as HR or marketing), it should be very easy for managers to hoard talent because the manager has a high degree of control over worker visibility. On the other hand, in areas such as engineering where many workers file patents, it should be more difficult for managers to hoard. [^17] High-visibility functions are defined as engineering, project management, and quality management. Low-visibility functions are defined as administration, finance, safety and health, HR, marketing, it, logistics, and purchasing. See Appendix Section E.2.1 for additional details on how these proxies are constructed and distributed in my sample.
Managers who face stronger incentives to hoard talent are substantially more likely to engage in hoarding behavior. Table 2 reports results from OLS regressions of a binary indicator for hoarding-prone managers based on my measure of suppressing potential ratings on the three proxies for hoarding incentives. Column 1 shows that managers with a 1 percentage point higher share of performance-related compensation (i.e., stronger financial incentives for hoarding) are 0.19 percentage points more likely to hoard talent ($p$ = 0.000). This effect corresponds to a 13 percentage point difference in predicted hoarding rates between the 90th and 10th percentiles of the financial incentive distribution. Column 2 uses team size as a proxy for hoarding incentives, and finds that managers with smaller teams are significantly more likely to hoard: a one-person increase in team size is associated with a 1.3 percentage point reduction in hoarding behavior ($p$ = 0.000), implying a 13 percentage point difference between the 90th and 10th percentiles of team size. Column 3 shows that managers in low-visibility functional areas are 4.0 percentage points more likely to hoard ($p$ = 0.002), relative to peers in high-visibility areas.
These results confirm that talent hoarding is substantially more common among managers who face stronger incentives to retain talent within their teams, in line with the predictions of my conceptual framework. Importantly, these findings are robust to using alternative specifications of the incentive measures, such as binary indicators for a high share of pay based on team performance or an indicator for leading a small team, as well as to alternative versions of the outcome variable, including the continuous measure of hoarding behavior (Appendix Tables  E4 and E5).
Higher levels of talent hoarding among managers with stronger incentives are also evident when analyzing a complementary measure of talent hoarding based on worker assignments to high-visibility trainings. At the firm, managers can assign their workers to in-person trainings, which are organized by a centralized and well-recognized department at the firm. These trainings are prestigious and typically increase worker visibility outside their team. For example, lists of training participants are widely circulated within the firm, and many workers list trainings on their CVs. I use the frequency with which managers restrict access to these training opportunities relative to what would be expected given worker characteristics to construct a measure of talent hoarding. The construction of this measure is analogous to that of the baseline measure using potential ratings. See Appendix Section E.2.3 for more details.
The effects of hoarding incentives on restricting access to high-visibility trainings are similar in both direction and economic magnitude to those observed using the potential ratings–based measure (Appendix Table B4). Managers whose performance-related compensation is 1 percentage point higher are 0.20 percentage points more likely to hoard talent ($p$ = 0.000). Similarly, a one-person increase in team size results in a 1.4 percentage point reduction in hoarding ($p$ = 0.000). Lastly, managers in low-visibility functional areas are 2.98 percentage points more likely to hoard ($p$ = 0.021). These patterns are highly robust to using alternative measures of both outcome and regressors (Appendix Tables E6 and E7).
In addition to the administrative analysis, which captures high-stakes realized outcomes, I use data from the manager survey to examine whether those facing stronger incentives are more likely to report talent hoarding. The advantage of the survey data is that the outcome measures are based on questions that explicitly elicit hoarding behavior, providing more direct insights into how managers perceive hoarding actions. I construct the incentive measures in the survey using a similar approach as in the administrative data (see Appendix Section E.2.4 for details), and analyze the two separate survey questions related to hoarding behavior described in Section 4.1.
Table 3 shows that the effects of hoarding incentives on perceptions of talent hoarding in the survey are similar in both direction and economic magnitude to those observed using administrative measures. Columns 1 to 3 use my primary question, which asks respondents how often managers feel the need to dissuade team members from exploring opportunities in other departments (i.e., perceived hoarding). See Appendix Table D10 for the raw response distribution and question wording. A 1 percentage point increase in the share of performance-related compensation leads to a 0.27 percentage point increase in perceived hoarding ($p$ = 0.019). Similarly, a one-person increase in team size results in a 0.4 percentage point decrease in perceived hoarding ($p$ = 0.026), and managers in low-visibility functional areas are 6.7 percentage points more likely to perceive hoarding ($p$ = 0.277).
Columns 4 to 6 use an alternative measure of perceived hoarding based on a question asking whether the risk of losing talent prevents managers from investing time and effort in their employees’ career development (see Appendix Table D7 for the raw response distribution and question wording). The results are similar: a 1 percentage point increase in performance compensation yields a 0.32 percentage point increase in perceived hoarding ($p$ = 0.038); a one-person increase in team size reduces perceived hoarding by 0.65 percentage points ($p$ = 0.003); and managers in low-visibility areas are 14.4 percentage points more likely to perceive hoarding ($p$ = 0.010).
These patterns are consistent across both survey questions and align closely with those in the administrative data, reinforcing the interpretation that stronger managerial incentives are associated with greater hoarding behavior. These findings are also robust to using alternative specifications of the incentive variables (Appendix Table E9). Placebo tests further confirm that these relationships are specific to talent hoarding and do not simply capture other organizational barriers (Appendix Table E10).
Taken together, these results provide evidence using several complementary measures of manager actions that managers are more likely to engage in talent hoarding if they face greater hoarding incentives. Interpreted through the lens of my conceptual framework, these results confirm that more misaligned incentives lead to higher levels of talent hoarding.